OpenStudy (anonymous):

Calculus I Problem: The table shows the position of a cyclist. t (seconds) 0 1 2 3 4 5 s (meters) 0 1.4 5.2 10.6 17.7 26.2 --> Estimate the instantaneous velocity when t = 3.

OpenStudy (anonymous):

What you need to do, @suzumiisama is you need to find the average velocity between t=2 and 3, find the average velocity between t=3 and 4 and average that. In other words, assuming this is a function \(s(t)\), your equation would be something like this: \[\frac{\frac{s(3)-s(2)}{3-2}+\frac{s(4)-s(3)}{4-3}}{2}\]

OpenStudy (anonymous):

Yay i got it :) Thanks so much

OpenStudy (anonymous):

My pleasure :)